OpenStudy (anonymous):

y = sqrt. of (36 - x^2)

OpenStudy (amistre64):

y = sqrt(36-x^2) square both sides: to get y^2 = 36 -x^2 x^2 + y^2 = 36 is the equation of a circle with center at origin and a radius of 6.

OpenStudy (amistre64):

the original equation is just for the top part of the circle